OpenStudy (anonymous):

What is the solution set of the inequality |3 - 2x| ≥ 4?

OpenStudy (anonymous):

|3 - 2x| ≥ 4 \[-4 \le 3 - 2x \le 4\]

OpenStudy (anonymous):

\[-4 - 3 \le -2x \le 4 - 3 \] \[-7 \le -2x \le 1\]

OpenStudy (anonymous):

\[-1 \le 2x \le 7\] \[-\frac{ 1 }{ 2 } \le x \le \frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

@KKJ |3 - 2x| ≥ 4 is NOT −4 ≤ 3−2x ≤ 4 http://www.wolframalpha.com/input/?i=%7C3+-+2x%7C+%E2%89%A5+4+

OpenStudy (anonymous):

It's actually -4 ≥ 3 - 2x and 3 - 2x ≥ 4